Hi can you help me with the following problem?
Write the answer in standard form of scientific notation correctly rounded to 2 figures.
(9.79 x 10^2) ( 4.06 x 10^-7)
______________________________
(8.33 x 10^-7) ( 2.59 x 10^4)
Thanks.
:).

Answer by checkley71(8403)

9.79*10^2*4.06*10^-7=39.75*10^-5 or 3.975*10^-4 numerator
8.33*10^-7*2.59*10^4=21.57*10^-3 or 2.157*10^-2 denominator
(3.975*10^-4)/(2.157*10^-2)
1.84*10^-2 or 1/1.84*10^2 answer.
1.84*10^-2 or 1/1.84*10^2